OpenStudy (anonymous):

Use mathematical induction to prove that the statement is true for every positive integer n. 2 is a factor of n^2 - n + 2..

OpenStudy (campbell_st):

start by proving its true for n = 1 the assume its true for n = k

OpenStudy (anonymous):

if n=1 then n2−n+2=12−1+2=2 which is a multiple of 2..

OpenStudy (anonymous):

now let's assume that n>1 and 2∣(n2−n+2) we have to prove that 2∣[(n+1)2−(n+1)+2]

OpenStudy (anonymous):

(n+1)2−(n+1)+2=n2+2n+1−n−1+2 =(n2−n+2)+2n=2⋅α+2n=2(α+n) and yeah don't forget that.... 2∣(n2−n+2)⇔n2−n+2=2⋅α this the induction hipothesis.

OpenStudy (campbell_st):

then in assuming its true for n = k you can say that n^2 - n + 2 = 2p now prove it for n = k + 1 so you have (k + 1)^2 - (k + 1) + 2 = k^2 + 2k + 1 - k - 1 + 2 = k^2 + k + 2 = k^2 - k + 2 + 2k using the assumption for 2p = 2p + 2k = 2(p + k) which is divisible by 2

OpenStudy (anonymous):

ohk...that's all, thanks for the help

OpenStudy (campbell_st):

oops should read for n = k k^2 - k + 2 = 2p

OpenStudy (anonymous):

@skylark06 r u frm india

OpenStudy (anonymous):

yeah!!

OpenStudy (anonymous):

which state in india

OpenStudy (anonymous):

@skylark06 u dereee